# SemanticSplat: Feed-Forward 3D Scene Understanding with Language-Aware Gaussian Fields

### v1

@@ Abstract.
Scene Understanding, 2D Foundational Model, Feature Distillation, Open-Vocabulary Scene Segmentation, Multi-Modal Representation Learning
submissionid: 1714ccs: Computing methodologies Scene understanding
Figure 1. Our approach utilizes sparse view images as input to reconstruct a holistic semantic Gaussian field, which includes both the Gaussian field with language features and the segmentation features. This reconstruction captures geometry, appearance, and multi-modal semantics, enabling us to perform multiple tasks such as novel view synthesis, depth prediction, open-vocabulary segmentation, and promptable segmentation.
1. Introduction
@@
Recently, 3D scene understanding methods leveraging powerful pre‑trained 2D foundational models—such as the Segment Anything Model (SAM) (Kirillov et al., 2023) and CLIP (Radford et al., 2021)—has emerged as a paradigm to enrich 3D representations with semantic knowledge distilled from readily available 2D data. However, directly transferring 2D semantic knowledge to 3D is non-trivial: 2D predictions often suffer from view-dependent inconsistencies, leading to noisy and unreliable semantic fields when aggregated across views. Besides, while NeRF (Mildenhall et al., 2020) and explicit 3D Gaussian splatting (Kerbl et al., 2023) methods enhanced with 2D features have shown promise for open‑vocabulary flexibility, existing approaches predominantly depend on per-scene optimization, making them impractical for dynamic or large‑scale applications.
@@
2. Related Work
2.1. 3D Scene Understanding
@@ 2. Related Work
Early language-aware scene representations embed CLIP features in NeRF volumes to support open-vocabulary queries, as demonstrated by LERF (Kerr et al., 2023a). Subsequent efforts migrate to 3D Gaussian Splatting (3DGS) for real-time rendering: GARField (Kim et al., 2024) distills SAM masks into Gaussians, LangSplat (Qin et al., 2024) auto-encodes a scene-wise language field, and Gaussian Grouping (Ye et al., 2025) attaches identity codes for instance-level clustering. Recent extensions such as SAGA (Cen et al., 2025) and 4D LangSplat (Li et al., 2025) provide promptable segmentation and temporally coherent language fields, respectively, yet their reliance on point-cloud surfaces limits mesh fidelity. Emerging approaches like OV-NeRF (Liao et al., 2024) introduce cross-view self-enhancement strategies to mitigate CLIP’s view inconsistency through semantic field distillation, while DiCo-NeRF (Choi et al., 2024) leverages CLIP similarity maps for dynamic object handling in driving scenes. Concurrent work MaskField (Gao et al., 2024) demonstrates how decomposing SAM mask features from CLIP semantics enables efficient 3D segmentation in Gaussian Splatting representation. Although fast in per-scene training, it still needs per-scene optimization.
2.2. Feed-Forward Gaussian Splatting
@@
Feed-forward reconstructors amortize 3DGS inference. For example, PixelSplat (Charatan et al., 2024) learns Gaussians from two views, while Splatter Image (Szymanowicz et al., 2024) accelerates single-view object recovery through per-pixel Gaussian prediction. The recent Hierarchical Splatter Image extension (Shen et al., 2024b) introduces parent-child Gaussian structures to recover occluded geometry through view-conditioned MLPs. To leverage multi-view cues, MVSplat (Chen et al., 2024c) builds cost volumes; we push this idea further by injecting monocular depth priors for texture-less scenes. Large-scale models trade hand-crafted geometry for data-driven priors: LGM (Tang et al., 2024), GRM (Xu et al., 2024), GS-LRM (Zhang et al., 2024), and LaRA (Chen et al., 2024b) reconstruct scenes in milliseconds but demand ¿60 GPU-days for pre-training. Gamba (Shen et al., 2024a) achieves 1000× speedup over optimization methods through Mamba-based sequential prediction of 3D Gaussians, though constrained to object-level reconstruction. Our approach reaches comparable quality in two GPU-days and, unlike LRMs, can be pre-trained with inexpensive posed images without depth supervision.
2.3. Lifting 2D Foundation Models to 3D
@@
Neural fields can aggregate multi-view image features into a canonical 3D space. Semantic NeRF (Zhi et al., 2021) and Panoptic Lifting (Siddiqui et al., 2023) fuse segmentation logits, showing that consistent 3D fusion cleans noisy 2D labels. Beyond labels, Distilled Feature Fields (Shen et al., 2023), LERF (Kerr et al., 2023b), NeRF-SOS (Fan et al., 2022), and FeatureNeRF (Ye et al., 2023) render pixel-aligned DINO or CLIP embeddings for tasks such as key-point transfer. Recent 3DGS adaptations (Lee et al., 2025; Liu et al., 2024; Qin et al., 2024; Shi et al., 2024; Li et al., 2024; Zhou et al., 2024a; Zuo et al., 2024; Zhou et al., 2025; Fan et al., 2024) adopt similar strategies to distill information from well-trained 2D models to 3D Gaussians. Feature 3DGS (Zhou et al., 2024a) generalizes distillation to explicit Gaussians; concurrent works like FMGS (Zuo et al., 2024) and SPLAT-Raj (Liu et al., 2024) confirm that SAM or LSeg signals can be attached to Gaussians for open-vocabulary editing.
3. Method
@@
Figure 2. We employ multiview transformers with cross-attention to extract features from multi-view images and use cost volumes for feature matching (see Sec. 3.1). Utilizing the multi-conditioned semantic features from Visual Feature Modules (VFMs) aggregated with the cost volumes (see Sec. 3.2), we predict Semantic Anisotropic Gaussians(see Sec. 3.3). Through a two-stage feature distillation process involving both segmentation(see Sec. 3.4) and language features(see Sec. 3.5), we reconstruct the holistic semantic feature field by jointly enforcing photometric fidelity and semantic consistency.
3.1. Efficient Depth Map Prediction
The first step of our SemanticSplat is to predict the depth map from the given inputs for further initiating the Gaussians. Typical image-to-image depth estimation pipelines like ViT-based encoder-decoder (Dosovitskiy et al., 2020) often lead to noisy edge artifacts in rendering results, caused by directly regressing point maps from image pairs. Therefore, we instead estimate depth maps for both target and source RGB images through feature matching with cost volumes. This method aggregates feature similarities across views, thereby enhancing the model’s cross-view awareness. These depth maps are then converted to point clouds, upon which Gaussian parameters are regressed.
@@ 3.1. Efficient Depth Map Prediction
We employ a CNN to extract down-sampled features from input views. These features are then processed by a Swin-Transformer (Liu et al., 2022; Xu et al., 2022, 2023), equipped with cross-attention layers to propagate information across views, enhancing the model’s ability to capture inter-view relationships. The resulting multi-view-aware features are represented as ${\left\{F_{i}\in{\mathbb{R}^{\frac{H}{s}\times{\frac{W}{s}}\times{C}}}\right\}_{i=1}^{N}}$, where $s$ is the down sampling factor and $C$ is the feature dimension. Cross-attention is applied bidirectionally across all views, generalizing the framework to arbitrary numbers of input images.
@@
Following MVSplat (Chen et al., 2024c), we adopt a plane-sweep stereo (Xu et al., 2023; Collins, 1996) approach for cross-view feature matching. For each view $i$ , we uniformly sample $D$ depth candidates $\left\{d_{m}\right\}_{m=1}^{D}$ from the near-to-far depth range. Features from another view $j$ are warped to view $i$ at each depth candidate $d_{m}$, generating $D$ warped features $\left\{F_{d_{m}}^{j\to{i}}\right\}_{m=1}^{D}$ The correlation between these warped features and view $i$ ’s original features is computed to construct the cost volume $\left\{C_{i}\in{\mathbb{R}^{\frac{H}{s}\times{\frac{W}{s}}\times{D}}}\right\}$. Finally, a 2D U-Net with a softmax layer predicts the per-view depth map by processing the concatenated Transformer features and cost volumes.
3.2. Multi-conditioned Semantic Features Aggregation
Recent Works (Chen et al., 2024a) investigate the geometry awareness and texture awareness of visual foundation models (VFMs) (Bommasani et al., 2021), which can enhance scene understanding. Leveraging the capabilities of VFMs, we aggregate pre-trained monocular multi-task semantic features into the cost volume to address challenging scenarios.
@@ 3.2. Multi-conditioned Semantic Features Aggregation
We leverage the pre-trained segmentation backbone from the Segment Anything Model (SAM) (Kirillov et al., 2023)and CLIP-LSeg model (Li et al., 2022) to get monocular features for each view. By interpolating to align with the cost volume resolution (Sec.3.1), the processed segmentation-semantic features ${\left\{F{{}_{i}^{SAM}}\in{\mathbb{R}^{\frac{H}{s}\times{\frac{W}{s}}\times{C_{SAM}}}}\right\}_{i=1}^{N}}$ from SAM and language-semantic features ${\left\{F{{}_{i}^{LSeg}}\in{\mathbb{R}^{\frac{H}{s}\times{\frac{W}{s}}\times{C_{LSeg}}}}\right\}_{i=1}^{N}}$ from CLIP-LSeg are concatenated with cost volumes$\left\{C_{i}\in{\mathbb{R}^{\frac{H}{s}\times{\frac{W}{s}}\times{D}}}\right\}_{i=1}^{N}$ and processed by a lightweight 2D U-Net (Rombach et al., 2022; Ronneberger et al., 2015) to regress a unified latent feature map, integrating geometric and semantic cues.
3.3. Semantic Anisotropic Gaussians Prediction
Despite significant progress in scene understanding and language-guided reconstruction, existing methods (Fan et al., 2024; Zhi et al., 2021; Qin et al., 2024; Li et al., 2024) often exhibit limited holistic scene comprehension and rely on single-modal segmentation. To address this, we propose holistic semantic field reconstruction via disentangled segmentation feature distillation and language feature distillation, implemented through anisotropic semantic Gaussians.
Compared to conventional Gaussians ${\left\{({\mu_{j},\alpha_{j},\sum_{j},c_{j}})\right\}_{j=1}^{n}}$ (Kerbl et al., 2023), semantic anisotropic Gaussians ${\left\{({\mu_{j},\alpha_{j},\sum_{j},c_{j},f_{j}})\right\}_{j=1}^{n}}$ incorporate latent space f into Gaussian attributes to represent 3D semantic fields, enabling joint rendering of novel-view RGB map C and semantic feature map F.
(1) $\begin{cases}\textbf{C}=\sum_{i=1}^{n}\textbf{c}_{i}\alpha_{i}G_{i}(X)\prod_{i=1}^{i-1}(1-\alpha_{j}G_{j}(X)),\\ \textbf{F}=\sum_{i=1}^{n}\textbf{f}_{i}\alpha_{i}G_{i}(X)\prod_{i=1}^{i-1}(1-\alpha_{j}G_{j}(X)),\end{cases}$
@@ 3.3. Semantic Anisotropic Gaussians Prediction
Pre-trained visual foundation models (VFMs) (Bommasani et al., 2021) often yield feature maps lacking view consistency and spatial awareness. To address this, we introduce a two-stage semantic feature distillation framework, including Segmentation Feature Field Distillation (Sec. 3.4) and Hierarchical-Context-Aware Language Field Distillation (Sec. 3.5) that lifts 2D features to 3D and refines latent feature maps, integrating diverse feature fields for holistic scene understanding.
3.4. Segmentation Feature Field
We leverage the Segment Anything Model (SAM) (Kirillov et al., 2023)—an advanced promptable segmentation model supporting inputs like points and bounding boxes—to distill segmentation-semantic embeddings into anisotropic Gaussians.
@@ 3.4. Segmentation Feature Field
(2) $L_{dist}^{Seg}=1-\texttt{sim}(f_{expand}(S),\hat{S})=1-\frac{S\cdot{\hat{S}}}{||S||\cdot{||\hat{S}||}}$
@@
We integrate SAM’s pre-trained mask decoder into our pipeline to generate segmentation masks. A consistency loss enforces alignment between masks derived from our image embeddings $M={\left\{M_{i}\in{\mathbb{R}^{H\times{W}}}\right\}_{i=1}^{M}}$ and SAM embeddings $\hat{M}={\left\{\hat{M}_{i}\in{\mathbb{R}^{H\times{W}}}\right\}_{i=1}^{M}}$, ensuring promptable segmentation compatibility. We employ a linear combination of Focal Loss (Lin et al., 2017) and Dice Loss (Milletari et al., 2016) in a 20:1 ratio for optimization.
@@
(4) $d=1-\frac{2\left|X\cap Y\right|}{\left|X\right|+\left|Y\right|}$
(5) $L_{mask}^{Seg}=FL(p_{t})+\frac{1}{20}d$
3.5. Hierarchical-Context-Aware Language Field
@@
With the segmentation feature branch (Sec. 3.4) frozen for stability, we use a new language-semantic head to predict the language-semantic features $f_{j}^{L}$ from the Gaussian semantic latent attribute $f_{j}$, as show in Figure 2, and then rasterize it into 2D language feature maps $L={\left\{L_{i}\in{\mathbb{R}^{h^{{}^{\prime\prime}}\times{w^{{}^{\prime\prime}}}\times{d^{{}^{\prime\prime}}}}}\right\}_{i=1}^{N}}$ and expand it to the dimension of CLIP-LSeg (Li et al., 2022) feature maps $\hat{L}={\left\{\hat{L_{i}}\in{\mathbb{R}^{H^{{}^{\prime\prime}}\times{W^{{}^{\prime\prime}}}\times{C^{{}^{\prime\prime}}}}}\right\}_{i=1}^{N}}$ as $\overline{L}$ for loss computation.
(6) $L_{dist}^{Lang}=1-\texttt{sim}(\overline{L},\hat{L})=1-\frac{\overline{L}\cdot{\hat{L}}}{||\overline{L}||\cdot{||\hat{L}||}}$
@@
(7) $\overline{L}^{h}=\frac{\sum\overline{L}\cdot{M^{h}}}{\sum{M^{h}}}$
3.6. Training Loss

### v2

@@ Abstract
Figure 1: Our approach utilizes sparse view images as input to reconstruct a holistic semantic Gaussian field, which includes both the Gaussian field with language features and the segmentation features. This reconstruction captures geometry, appearance, and multi-modal semantics, enabling us to perform multiple tasks such as novel view synthesis, depth prediction, open-vocabulary segmentation, and promptable segmentation.
1 Introduction
@@
Recently, 3D scene understanding methods leveraging powerful pre‑trained 2D foundational models—such as the Segment Anything Model (SAM) [20] and CLIP [33]—has emerged as a paradigm to enrich 3D representations with semantic knowledge distilled from readily available 2D data. However, directly transferring 2D semantic knowledge to 3D is non-trivial: 2D predictions often suffer from view-dependent inconsistencies, leading to noisy and unreliable semantic fields when aggregated across views. Besides, while NeRF [29] and explicit 3D Gaussian splatting [15] methods enhanced with 2D features have shown promise for open‑vocabulary flexibility, existing approaches predominantly depend on per-scene optimization, making them impractical for dynamic or large‑scale applications.
@@
2 Related Work
2.1 3D Scene Understanding
@@ 2 Related Work
Early language-aware scene representations embed CLIP features in NeRF volumes to support open-vocabulary queries, as demonstrated by LERF [16]. Subsequent efforts migrate to 3D Gaussian Splatting (3DGS) for real-time rendering: GARField [18] distills SAM masks into Gaussians, LangSplat [32] auto-encodes a scene-wise language field, and Gaussian Grouping [48] attaches identity codes for instance-level clustering. Recent extensions such as SAGA [2] and 4D LangSplat [24] provide promptable segmentation and temporally coherent language fields, respectively, yet their reliance on point-cloud surfaces limits mesh fidelity. Emerging approaches like OV-NeRF [25] introduce cross-view self-enhancement strategies to mitigate CLIP’s view inconsistency through semantic field distillation, while DiCo-NeRF [7] leverages CLIP similarity maps for dynamic object handling in driving scenes. Concurrent work MaskField [13] demonstrates how decomposing SAM mask features from CLIP semantics enables efficient 3D segmentation in Gaussian Splatting representation. Although fast in per-scene training, it still needs per-scene optimization.
2.2 Feed-Forward Gaussian Splatting
@@
Feed-forward reconstructors amortize 3DGS inference. For example, PixelSplat [3] learns Gaussians from two views, while Splatter Image [41] accelerates single-view object recovery through per-pixel Gaussian prediction. The recent Hierarchical Splatter Image extension [36] introduces parent-child Gaussian structures to recover occluded geometry through view-conditioned MLPs. To leverage multi-view cues, MVSplat [6] builds cost volumes; we push this idea further by injecting monocular depth priors for texture-less scenes. Large-scale models trade hand-crafted geometry for data-driven priors: LGM [42], GRM [46], GS-LRM [49], and LaRA [4] reconstruct scenes in milliseconds but demand ¿60 GPU-days for pre-training. Gamba [37] achieves 1000× speedup over optimization methods through Mamba-based sequential prediction of 3D Gaussians, though constrained to object-level reconstruction. Our approach reaches comparable quality in two GPU-days and, unlike LRMs, can be pre-trained with inexpensive posed images without depth supervision.
2.3 Lifting 2D Foundation Models to 3D
@@
Neural fields can aggregate multi-view image features into a canonical 3D space. Semantic NeRF [51] and Panoptic Lifting [40] fuse segmentation logits, showing that consistent 3D fusion cleans noisy 2D labels. Beyond labels, Distilled Feature Fields [38], LERF [17], NeRF-SOS [11], and FeatureNeRF [47] render pixel-aligned DINO or CLIP embeddings for tasks such as key-point transfer. Recent 3DGS adaptations [21, 27, 32, 39, 23, 53, 56, 54, 12] adopt similar strategies to distill information from well-trained 2D models to 3D Gaussians. Feature 3DGS [53] generalizes distillation to explicit Gaussians; concurrent works like FMGS [56] and SPLAT-Raj [27] confirm that SAM or LSeg signals can be attached to Gaussians for open-vocabulary editing.
3 Method
@@
Figure 2: We employ multiview transformers with cross-attention to extract features from multi-view images and use cost volumes for feature matching (see Sec. 3.1). Utilizing the multi-conditioned semantic features from Visual Feature Modules (VFMs) aggregated with the cost volumes (see Sec. 3.2), we predict Semantic Anisotropic Gaussians(see Sec. 3.3). Through a two-stage feature distillation process involving both segmentation(see Sec. 3.4) and language features(see Sec. 3.5), we reconstruct the holistic semantic feature field by jointly enforcing photometric fidelity and semantic consistency.
3.1 Efficient Depth Map Prediction
The first step of our SemanticSplat is to predict the depth map from the given inputs for further initiating the Gaussians. Typical image-to-image depth estimation pipelines like ViT-based encoder-decoder [10] often lead to noisy edge artifacts in rendering results, caused by directly regressing point maps from image pairs. Therefore, we instead estimate depth maps for both target and source RGB images through feature matching with cost volumes. This method aggregates feature similarities across views, thereby enhancing the model’s cross-view awareness. These depth maps are then converted to point clouds, upon which Gaussian parameters are regressed.
@@ 3.1 Efficient Depth Map Prediction
We employ a CNN to extract down-sampled features from input views. These features are then processed by a Swin-Transformer [28, 44, 45], equipped with cross-attention layers to propagate information across views, enhancing the model’s ability to capture inter-view relationships. The resulting multi-view-aware features are represented as ${\left\{F_{i}\in{\mathbb{R}^{\frac{H}{s}\times{\frac{W}{s}}\times{C}}}\right\}_{i=1}^{N}}$, where $s$ is the down sampling factor and $C$ is the feature dimension. Cross-attention is applied bidirectionally across all views, generalizing the framework to arbitrary numbers of input images.
@@
Following MVSplat [6], we adopt a plane-sweep stereo [45, 8] approach for cross-view feature matching. For each view $i$ , we uniformly sample $D$ depth candidates $\left\{d_{m}\right\}_{m=1}^{D}$ from the near-to-far depth range. Features from another view $j$ are warped to view $i$ at each depth candidate $d_{m}$, generating $D$ warped features $\left\{F_{d_{m}}^{j\to{i}}\right\}_{m=1}^{D}$ The correlation between these warped features and view $i$ ’s original features is computed to construct the cost volume $\left\{C_{i}\in{\mathbb{R}^{\frac{H}{s}\times{\frac{W}{s}}\times{D}}}\right\}$. Finally, a 2D U-Net with a softmax layer predicts the per-view depth map by processing the concatenated Transformer features and cost volumes.
3.2 Multi-cond Semantic Features Aggregation
Recent Works [5] investigate the geometry awareness and texture awareness of visual foundation models (VFMs) [1], which can enhance scene understanding. Leveraging the capabilities of VFMs, we aggregate pre-trained monocular multi-task semantic features into the cost volume to address challenging scenarios.
@@ 3.2 Multi-cond Semantic Features Aggregation
We leverage the pre-trained segmentation backbone from the Segment Anything Model (SAM) [20]and CLIP-LSeg model [22] to get monocular features for each view. By interpolating to align with the cost volume resolution (Sec.3.1), the processed segmentation-semantic features ${\left\{F{{}_{i}^{SAM}}\in{\mathbb{R}^{\frac{H}{s}\times{\frac{W}{s}}\times{C_{SAM}}}}\right\}_{i=1}^{N}}$ from SAM and language-semantic features ${\left\{F{{}_{i}^{LSeg}}\in{\mathbb{R}^{\frac{H}{s}\times{\frac{W}{s}}\times{C_{LSeg}}}}\right\}_{i=1}^{N}}$ from CLIP-LSeg are concatenated with cost volumes$\left\{C_{i}\in{\mathbb{R}^{\frac{H}{s}\times{\frac{W}{s}}\times{D}}}\right\}_{i=1}^{N}$ and processed by a lightweight 2D U-Net [34, 35] to regress a unified latent feature map, integrating geometric and semantic cues.
3.3 Semantic Anisotropic Gaussians Prediction
Despite significant progress in scene understanding and language-guided reconstruction, existing methods [12, 51, 32, 23] often exhibit limited holistic scene comprehension and rely on single-modal segmentation. To address this, we propose holistic semantic field reconstruction via disentangled segmentation feature distillation and language feature distillation, implemented through anisotropic semantic Gaussians.
Compared to conventional Gaussians ${\left\{({\mu_{j},\alpha_{j},\sum_{j},c_{j}})\right\}_{j=1}^{n}}$ [15], semantic anisotropic Gaussians ${\left\{({\mu_{j},\alpha_{j},\sum_{j},c_{j},f_{j}})\right\}_{j=1}^{n}}$ incorporate latent space f into Gaussian attributes to represent 3D semantic fields, enabling joint rendering of novel-view RGB map C and semantic feature map F.
$\begin{cases}\textbf{C}=\sum_{i=1}^{n}\textbf{c}_{i}\alpha_{i}G_{i}(X)\prod_{i=1}^{i-1}(1-\alpha_{j}G_{j}(X)),\\ \textbf{F}=\sum_{i=1}^{n}\textbf{f}_{i}\alpha_{i}G_{i}(X)\prod_{i=1}^{i-1}(1-\alpha_{j}G_{j}(X)),\end{cases}$ (1)
@@ 3.3 Semantic Anisotropic Gaussians Prediction
Pre-trained visual foundation models (VFMs) [1] often yield feature maps lacking view consistency and spatial awareness. To address this, we introduce a two-stage semantic feature distillation framework, including Segmentation Feature Field Distillation (Sec. 3.4) and Hierarchical-Context-Aware Language Field Distillation (Sec. 3.5) that lifts 2D features to 3D and refines latent feature maps, integrating diverse feature fields for holistic scene understanding.
3.4 Segmentation Feature Field
We leverage the Segment Anything Model (SAM) [20]—an advanced promptable segmentation model supporting inputs like points and bounding boxes—to distill segmentation-semantic embeddings into anisotropic Gaussians.
@@ 3.4 Segmentation Feature Field
$L_{dist}^{Seg}=1-\texttt{sim}(f_{expand}(S),\hat{S})=1-\frac{S\cdot{\hat{S}}}{||S||\cdot{||\hat{S}||}}$ (2)
@@
We integrate SAM’s pre-trained mask decoder into our pipeline to generate segmentation masks. A consistency loss enforces alignment between masks derived from our image embeddings $M={\left\{M_{i}\in{\mathbb{R}^{H\times{W}}}\right\}_{i=1}^{M}}$ and SAM embeddings $\hat{M}={\left\{\hat{M}_{i}\in{\mathbb{R}^{H\times{W}}}\right\}_{i=1}^{M}}$, ensuring promptable segmentation compatibility. We employ a linear combination of Focal Loss [26] and Dice Loss [30] in a 20:1 ratio for optimization.
@@
$d=1-\frac{2\left|X\cap Y\right|}{\left|X\right|+\left|Y\right|}$ (4)
$L_{mask}^{Seg}=FL(p_{t})+\frac{1}{20}d$ (5)
3.5 Hierarchical-Context-Aware Language Field
@@
With the segmentation feature branch (Sec. 3.4) frozen for stability, we use a new language-semantic head to predict the language-semantic features $f_{j}^{L}$ from the Gaussian semantic latent attribute $f_{j}$, as show in Figure 2, and then rasterize it into 2D language feature maps $L={\left\{L_{i}\in{\mathbb{R}^{h^{{}^{\prime\prime}}\times{w^{{}^{\prime\prime}}}\times{d^{{}^{\prime\prime}}}}}\right\}_{i=1}^{N}}$ and expand it to the dimension of CLIP-LSeg [22] feature maps $\hat{L}={\left\{\hat{L_{i}}\in{\mathbb{R}^{H^{{}^{\prime\prime}}\times{W^{{}^{\prime\prime}}}\times{C^{{}^{\prime\prime}}}}}\right\}_{i=1}^{N}}$ as $\overline{L}$ for loss computation.
$L_{dist}^{Lang}=1-\texttt{sim}(\overline{L},\hat{L})=1-\frac{\overline{L}\cdot{\hat{L}}}{||\overline{L}||\cdot{||\hat{L}||}}$ (6)
@@
$\overline{L}^{h}=\frac{\sum\overline{L}\cdot{M^{h}}}{\sum{M^{h}}}$ (7)
3.6 Training Loss